OpenStudy (anonymous):

Linda has two porcelain spheres. The radii of the spheres are in a ratio of 2 : 3. What is the ratio between the volumes of the two spheres? Answer 8 : 27 8 : 81 16 : 27 16 : 81

OpenStudy (anonymous):

radii are measured in linear units, like ft or cm, but volumes are measured in cubic units, like ft^3 or in^3. 2^3=? 3^3=?

OpenStudy (anonymous):

2^3 = 6 and 3^3= 9?

OpenStudy (anonymous):

My bad, im not asking what is 2*3, that would be 6\[2^3=2\cdot 2\cdot 2\]\[3^3=3\cdot 3\cdot 3\]Thats what i'm looking for.

OpenStudy (anonymous):

8 27

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome :)